OpenStudy (anonymous):

Find the exact arc length of the portion of the graph of y=(1/6)x^3+1/(2x) from x=1 to x=2.

OpenStudy (anonymous):

I have f'(x)=(x^2)/2 - 1/(2x^2) so \[\int\limits_{1}^{2} \sqrt{1+((x^4)/2)-(1/(4x^4))}dx\]